OpenStudy (anonymous):

Solve the following equations 3x - 4y =23 -7x - 6y = 23

OpenStudy (tkhunny):

Have you considered affecting a solution?

OpenStudy (anonymous):

I know you have to isolate the variable on one equation, I just don't know how to solve exactly.

OpenStudy (tkhunny):

Can you solve for x? 3x - 4y = 23

OpenStudy (anonymous):

I don't know how @tkhunny

OpenStudy (tkhunny):

I don't believe that. Add "4y" to both sides and see what happens.

OpenStudy (anonymous):

3x - 8y = 27y??

OpenStudy (tkhunny):

-4y + 4y = 0 -- Why did you subtract 4y? 23 + 4y = 23 + 4y -- Add together only "LIKE" terms. You do not appear to have the algebra skills to solve this problem. Let's test. What is an Additive Inverse? Provide an example. How do we identify "Like" terms. Provide at least two examples.